Update(9:30AM): All suspect hardware has been replaced. We believe affected customers should now be operational and will continue to monitor the situation.
Update(7:15AM): We are still working towards resolution with the remaining customers that are down
Update(3:03AM): We were able to restore most subscribers via equipment reboot, however a subset of circuits cannot be restored without hardware replacement. We are dispatching with replacement hardware with an ETA of 5AM.
At approximately 10:40PM, equipment serving a subset of legacy DSL subscribers in Northern California had a hardware failure. We are working to restore traffic as quickly as possible but anticipate resolution will require an emergency maintenance reload of the device. Updates to follow.
– Network Engineering